OpenStudy (beastie234):

An object is falling from a height of 7.5 meters. At what height will its velocity be 7 meters/second? a. 49 meters b. 7.5 meters c. 7.0 meters d. 5.0 meters e. 2.5 meters

OpenStudy (sharmagk83):

Data Given h = 7.5 m, v = 7 m/s and g = 9.81 m/s From third equation of motion we know that \[v^{2} - u^{2} = 2as\]

OpenStudy (sharmagk83):

\[7^{2}- 0^{2} = 2\times9.81\times h_{1}\] \[h_{1} = \frac{ (7)^{2} }{ 2\times9.81 } = 2.5 m\] Height at which velocity is 7 m/s is \[h' = h-h_{1} = 7.5 m - 2.5 m = 5.0 m\] So Option D is correct.